Convert -5pi/2 radians into degrees (report answer between 0 and 360 degrees)
Thanks,
paulaficarella@gmail.com

Answer by Alan3354(69443)

Convert -5pi/2 radians into degrees (report answer between 0 and 360 degrees)
----------
-5pi/2*(180/pi) = -450 degs
Add 360 repeatedly until you get a + number
150 + 360 = -90
-90 + 360 = 270 degs